OpenStudy (unklerhaukus):

\[(x−2y+1)dx+(4x−3y−6)dy=0\]

OpenStudy (unklerhaukus):

\[\frac{\text dy}{\text dx}=\frac{2y-x-1}{4x-3y-6}\]

OpenStudy (experimentx):

\[ \frac{dy}{dx} = \frac{(x−2y+1)}{(4x−3y−6)}\] Let, X = x+h, and Y = y+k \[ \frac{dY}{dX} = \frac{(X−2Y)+ (h - 2k + 1)}{((4X−3Y)+(4h-3y−6)}\] Now, h - 2k + 1 = 0 4h - 3k - 6 = 0 http://www.wolframalpha.com/input/?i=solve++h+-+2k+%2B+1+%3D+0%2C+4h+-+3k+-+6+%3D+0 Solve the above homogeneous differential equation above in X and Y, and put the values X and Y back to x and y from h and k

OpenStudy (unklerhaukus):

your missing a negative on your first line

OpenStudy (experimentx):

Oh .. sorry about that!!

OpenStudy (anonymous):

for this kind of questions, first find the solution for the equations in numerator and denominator then shift the origin to that point now use y=zx it becomes easier to solve

OpenStudy (anonymous):

@experimentX 's method is right :)

OpenStudy (experimentx):

reducible to homogeneous form!!!

OpenStudy (unklerhaukus):

\[\frac{\text d X}{ \text d Y}=\frac{2Y+2k-X-h-1}{4X+4h-3Y-3k-6}\]\[=\frac{2Y-X+(2k-h-1)}{4X-3Y+(4h-3k-6)}\]\[(2k-h-1)=0;\qquad(4h-3k-6)=0\]

OpenStudy (experimentx):

\[ \frac{dY}{dX} = \frac{2Y - X}{4X - 3Y}\] let Y = vX <--- usual method

OpenStudy (unklerhaukus):

\[k=2\]\[h=3\]

OpenStudy (unklerhaukus):

\[\frac{\text d Y}{\text d X}=\frac{2Y-X}{4X-3Y}=\frac{2(Y/X)-1}{4-3(Y/X)}\] \[V=(Y/X);\qquad Y=VX\]\[\frac{\text d Y}{\text d X}=X\frac{\text d V}{\text d X}+V\]

OpenStudy (unklerhaukus):

\[X\frac{\text d V}{\text d X}=\frac{2V-1}{4-3V}-V=\frac{2V-1}{4-3V}-V\frac{4-3V}{4-3V}\]\[X\frac{\text d V}{\text d X}=\frac{3V^2-2V-1}{4-3V}\]

OpenStudy (unklerhaukus):

\[\frac{\text d X}{X}=\frac{4-3V}{3V^2-2V-1}\text d V \]

OpenStudy (unklerhaukus):

what is the next step , unless i have errors so far

OpenStudy (experimentx):

\[ \int \frac{1}{X} dX = \ln X\]

OpenStudy (experimentx):

I think we can solve the other guy using partial fraction http://www.wolframalpha.com/input/?i=partial+fraction+%284-3x%29%2F%283x^2+-+2x+-+1%29

OpenStudy (unklerhaukus):

\[\ln X=\int \frac{1}{4(V-1)}-\frac{15}{4(3V+1)}\text d V \] \[=\frac{1}{4}\int\frac{1}{(V-1)}-\frac{15}{(3V+1)}\text d V\]

OpenStudy (experimentx):

\[ \frac{1}{4} \ln (v-1) - 5 \ln(2V + 1)\]

OpenStudy (unklerhaukus):

2V or 3V?

OpenStudy (experimentx):

\[ \ln X = \frac{1}{4} \ln (V-1) - 5 \ln(3V + 1) + C\] Putting back values of V, \[ \ln X = \frac{1}{4} \ln (Y/X-1) - 5 \ln(3Y/X + 1) + C\] put values of X and Y X = x + 3 Y = y + 2 then you have answer.

OpenStudy (unklerhaukus):

\[4\ln(X+3)=\ln\left(\frac{y+2}{x+3}-1\right)-5\ln\left(3\frac{y+2}{x+3}+1\right)+c\] \[(X+3)^4={\frac{y+2}{x+3}-1}+\left(3\frac{y+2}{x+3}+1\right)^{-5} +e^c \]?

OpenStudy (unklerhaukus):

X*x

OpenStudy (unklerhaukus):

\[(x+3)^4={\frac{y+2}{x+3}-1}+\left(3\frac{y+2}{x+3}+1\right)^{-5} +e^c \]is that the answer to my question/

OpenStudy (unklerhaukus):

ah bother